2. The electric potential at the exact center of a square is 36,000 V when there are charges of 3.0 μC in each of the four corners of the square.

a. Calculate the distance from each comer of the square to the center of the square.

b. A proton is released from rest at the center of the above square. What is the speed of the proton when it is very far away from the square?
Answer #1

a) let the distance is d

charge = 4*k * Q/r

36000 = 9 *10^9 * 4 * 3 *10^-6/(d)

solving for d

d = 3 m

the distance from each corner is 3 m

b) let the speed of the proton is v

Using conservation of energy

0.50 * 1.67 *10^-27 * v^2 = 36000 * 1.602 *10^-19

solving for v

v= 2.63 *10^8 m/s

the speed of proton will be 2.63 *10^8 m/s